A cockroach of mass m lies on the rim of a uniform disk of mass 5.00 m that can rotate freely about its center like a merry-go-round. Initially the cockroach and disk rotate together with an angular velocity of 0.410 rad/s. Then the cockroach walks halfway to the center of the disk.
(a) What then is the angular velocity of the cockroach-disk system?
(b) What is the ratio K/K0 of the new kinetic energy of the system to its initial kinetic energy?
Answer #1

Apply the angular momentum conservation, 1,0 = lo Calculate 1, for the disc and cockroach system, 1o = (5m)r? + mr 1= 3(5m)rº(b) The change in angular velocity is, Ao=0-0. 14 13 - = @(0.2727) =0.27270, The ratio of kinetic energies is, K. (loo 2 =*=